A multiproduct company produced three different kinds of products (A, B, and C). The following table gives a breakdown of the price, variable costs, and the expected number of units to be sold

Price Per Unit ($) Variable Costs Per Unit ($) Expected Units Sold Expected Units Sold (%)
Product A 100 50 100 50%
Product B 50 30 60 30%
Product C 20 10 40 20%

The firm's total fixed cost is $6,600.

What is the total number of units at which the company breaks even on their fixed costs? What is the total number of units at which the company breaks even on their fixed costs?
Answer #1

Contribution margin = Selling price - Variable cost

Product A = 100 - 50 = $50

Total contribution margin = $50

Ratio = 50%

Contribution margin as per mix

= 50*50%

= $25

Product B

= 50 - 30 = $20

Ratio = 30%

Contribution margin as per mix

= 20*30%

= $6

Product C

= 20 - 10

= $10

Ratio = 20%

Contribution margin as per mix

= 10*20%

= $2

Total contribution margin

= 25 + 6 + 2

= $33

Fixed cost = $6,600

Breakeven units

= Fixed cost/ Contribution margin per unit

= 6,600/33

= 200 units.

Therefore the correct answer is 200 units.

If you find the answer helpful please upvote.